A 12V Car Battery has an internal resistance of 0.2ohms. What potential difference must be applied across its terminals by a battery charger in order that the charging current will be 5A?

Grade:10

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vishnu

In case of charging

potential difference= E + Ir = 12 + 5x0.2 = 13V Answer

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..